You and your little brother are shifting a large 58 kg rock by carrying it together on a 3.0 m long litter.
Your little brother is capable of exerting an upwards force of 100 N on one end of the litter.

How far should the rock be placed from you in order for you and your brother to carry the litter without it tipping? (in m to 2 s.f)
1
Expert's answer
2020-04-03T13:22:57-0400


On this picture "F_1" is the force you provide, "100 N" is the force of your brouther and "mg = 58*10 = 580N" is the weight of the stone. "h_1" and "h_2" are corresponding shoulders.

In order to prevent  litter move up or down, all three forces must be balanced:

"F_1 + 100 = 580 \Rightarrow F_1 = 480N"

In order to prevent the litter to tip the angular momentum with respect to the left point of revolution should be balanced:

"580\cdot h_1 = 100\cdot 3 \Rightarrow h_1 \approx0.52"


Answer. The rock should be placed at the distance 0.52 m from you.